# Chasing Anharmonicities in Polarization-Orientation Raman Spectra of Acene Crystals with Machine Learning

### v1

@@ I Introduction
In this paper, we provide an ab initio quality framework that can be used to reproduce, explain and provide quantitative insights into the type of anharmonic effects that can be directly probed by PO-Raman spectroscopy. As a case study, we focus on linear oligoacene molecular crystals, in particular naphthalene and anthracene. In order to include all orders of anharmonicity of the potential energy surface (PES), our approach is based on the well-established time-correlation function formalism for Raman signals [11, 12]. The standard approach of combining some flavor of ab initio molecular dynamics (AIMD)[13] with density functional perturbation theory (DFPT) [14] to obtain Raman spectra is especially challenging in this case, due to the vast range of experimental conditions one needs to explore, and the large cell sizes required to accurately describe the phonon manifold. Furthermore, to capture subtle variations in the angle-resolved Raman intensity, the spectra have to be thoroughly converged, and finite-size artifacts due to periodic boundary conditions need to be mitigated. Naturally, machine-learning (ML) methods can be exploited in order to make these tasks feasible, but they also need to be scrutinized for accuracy. An array of related ML methods have been previously discussed and benchmarked in the literature [15, 16, 17, 18], but the focus on temperature-dependent PO-signals proves to be a more stringent test of the necessary accuracy of the framework.
In this manuscript, we begin by introducing the theoretical background of Raman scattering, emphasizing the limitations of the harmonic approximation and the need for anharmonic treatments. To address these challenges, we train machine-learning-based models for interatomic potentials and polarizability tensors, benchmarking their accuracy against reference ab initio data. The RGDOS approximation proposed by Komsa and coworkers in Refs. [19, 20] for the polarizability of large supercells is assessed in comparison to more stringent approaches. We then employ these models to simulate the temperature-dependent PO-Raman spectra of anthracene and naphthalene, capturing anharmonic effects related to temperature and the quantum nature of the nuclei. By comparing our results with experimental data, we analyze the polarization dependence of Raman intensities and discuss the subtle deviations from the harmonic predictions. Our findings highlight the interplay between anharmonic lattice dynamics and Raman tensor symmetries, providing insights into the origins of the temperature-dependent behavior of low-frequency Raman modes and suggesting guidelines for the interpretation of the rich amount of information contained in PO-dependent Raman spectra. This work demonstrates the power of combining machine learning with established theoretical frameworks to unravel complex anharmonic phenomena in molecular crystals.
@@ II Theory of Polarization-Orientation Raman Scattering
Single crystal polarization-orientation Raman experiments are often conducted in a back- (or forward-) scattering geometry, where the incident () and scattered () fields are linearly polarized and propagate through the crystal, which is positioned at a certain orientation [8, 22, 23]. The scattered radiation is then collected with polarization either perpendicular () or parallel () to that of the incident beam. Once the configuration is fixed, the experiment is repeated spanning an angle between and an arbitrary reference, either by rotating the sample or the direction of .
@@
where is the electronic ground state, the summation runs over all the excited electronic states, is the electric dipole operator and denote Cartesian components. With this definition of the polarizability tensor, Eq. 1 becomes the well known Placzek approximation [24, 25] to the full expression of the light scattering cross section, first derived by Kramers, Heisenberg and Dirac [26, 27]. Note that, as recently highlighted in Ref. [28], by treating the quantity in Eq. 3 as the polarizability tensor (or dielectric susceptibility) entering Eq. 1, we are already excluding the effect of finite scattering wave-vectors, essentially assuming that only vibrations at the -point of the crystal phonon Brillouin zone are probed in the experiment, i.e. , where the are the propagation directions of incoming and scattered fields. This tends to be a very good assumption for first-order Raman scattering, but breaks down in the context of second-order Raman scattering in polar materials [28].
@@
that will depend trivially on temperature through the population of the initial states. These derivations can also be found in the SI of Ref. [10]. We will normally refer to the angular dependence, for a given , as the “PO-pattern” of the corresponding mode in the following.
Eq. 7 is the well known expression for the (harmonic) polarized Raman scattering cross section that is commonly employed to rationalize experimental results in terms of Raman tensors and factor group analysis [30]. As shown above and already clearly stressed in [10], this simple expression is the result of what is often termed a “double harmonic” approximation, as it relies on both harmonic normal modes and the truncation at first order of the expansion in Eq. 4. Clearly, this formula would fall short of accounting for any effect of anharmonic lattice dynamics in the Raman spectrum. Particularly relevant for this work is the fact that in this approximation, the dependence of the (normalized) scattered intensity on the polarization angle is fully determined by the shape, symmetry and relative values of the Raman tensor components of a given mode, and by the geometry of the experiment. Notably, the thermal populations enter only as a multiplicative factor. As the space group of the crystal and the molecular sites do not change with temperature (assuming no phase transitions), the shape and symmetry of are fixed. A common interpretation of this observation is that besides a change in its absolute values, the angular dependence of the intensity calculated with Eq. 7 will be temperature independent in the harmonic picture. As we will see in this paper, though, keeping to Eq. 7 but including temperature-dependent lattice variations can lead to a change in sign and relative magnitude of certain components of , which in turn lead to pronounced variations in the PO pattern that are not connected to dynamical anharmonicity of the vibrational modes.
Any further change with temperature of the PO-patterns must be investigated taking into account dynamical anharmonicity of molecular vibrations, which presents an added challenge to simulations. In particular, dynamical methods which are not based on perturbation theory for the nuclear degrees of freedom, provide an attractive platform to include these effects in their totality. For practical purposes, it is convenient to switch from the energy representation of Eq. 1 to the Heisenberg representation in the time domain, resulting in
@@ III Simulating PO-Raman Signals
Achieving first-principles accuracy in this evaluation relies on being able to obtain calculations of forces and polarizability tensors, through the realization of AIMD trajectories and perturbation theory calculations, resulting in a very high computational cost [43]. Previous studies [1, 44, 45] have used AIMD to successfully compute anharmonic Raman spectra of molecular crystals, albeit computing very few trajectories of the order of 10 ps, resulting on a high uncertainty in peak intensities. In the present study, converged peak intensities are necessary in order to compare polarization-orientation patterns of different temperatures.
@@ III.1 Machine Learning Models for Polarizability Tensors
Both methods mentioned above have already been successfully used in the context of predicting polarizability tensors that were then used to calculate standard vibrational Raman spectra[15, 16, 49, 53, 48, 18]. Here we assess them for the more stringent test of predicting the PO-pattern of each Raman peak.
@@ III.2 The RGDOS Approximation to Polarizability Tensors
The RGDOS approximation to the polarizability tensor, as introduced in Ref. [19] and extended for molecular dynamics in Ref. [20], is an efficient approach to reconstruct the polarizability tensor of a large supercell as a weighted sum of polarizabilities of the multiple unit cells that constitute it. In Ref. [16], Berger et al. showed the good performance of this approximation in reproducing DFT-quality polarizabilities for large supercells and yielding good Raman spectra, including second-order and resonant contributions. The method relies on the existence of a unique equilibrium reference structure and its corresponding eigenmodes, being therefore unable to describe different phases of a material within a single model, but performing well when no phase transitions occur.
@@
where anharmonic effects are included in the correlation function of the projected supercell displacements. A further possible simplification regards the truncation of Eq. 13 with respect to the number of phonons considered. The evaluation of Eq. 14 then requires fewer Raman tensors and projection coefficients. We demonstrate in the SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results. (Section S2) how retaining only the intermolecular modes in the expansion has no effect in the resulting low-frequency spectrum of the polyacene crystals studied here, making this method particularly attractive for the investigation of low THz regions of molecular crystals.
@@ IV Results and Discussion
IV.1 Performance of MLIPs and Polarizability Models
In the following, we will compare different models for MLIPs and polarizability tensors of polyacene molecular crystals. Berger et al. have performed benchmarking of similar frameworks for different classes of inorganic materials[16]. Here, we apply this approach to the realm of molecular crystals and include a comparison with a state-of-the-art equivariant neural network model for polarizabilities, with the ultimate goal of obtaining ab initio quality PO-Raman spectra.
We start by assessing the quality of the MLIPs we trained for a crucial quantity in the simulation of Raman spectra, namely the frequencies of the crystal vibrations. A particularly challenging region, where we must guarantee that the errors are acceptable, is the low-THz part of the spectrum, which is determined by weak intermolecular interactions often poorly described in MLIPs that are inherently local. Even though we will employ these potentials in dynamical simulations, we choose the harmonic phonons at the -point of the Brillouin zone as a deterministic measure of the quality of our potentials, since they are free of any statistical uncertainties. As detailed in Methods Section VI.1, we compare a second-generation Behler-Parrinello high-dimensional neural network potential [54], which we label HDNNP-MLIP and a equivariant graph neural-network MACE [46] interatomic potential, which we label MACE-MLIP.
We take the anthracene molecular crystal as a benchmark case for evaluating the quality of the machine-learned models for the PES. Details of the training dataset and procedure are given in Section VI. As shown in Fig. 1a MACE-MLIP outperforms the HDNNP-MLIP in the prediction of -point harmonic phonon frequencies resulting in a lower average percentage error of 0.2% against the 1.6% obtained from HDNNP-MLIP. Moreover, MACE-MLIP exhibits a significantly tighter distribution of errors, with a low median and minimal variability, highlighting the consistency and accuracy of this model. In contrast, HDNNP-MLIP shows a broader error distribution, indicating less reliable predictions. When focusing specifically on the low THz region, the error on the harmonic frequency of Raman active modes reported in Fig. 1b is almost negligible for the MACE-MLIP model, while the HDNNP-MLIP largely underestimates the harmonic frequency of the three highest frequency librational modes, with errors up to 20%. These conclusions are consistent with the broader benchmark on accuracy of vibrational properties of ML potentials reported in Ref. [17].
Next, we assess the quality of our ML models for the prediction of the polarizability tensors. In Fig. 2 we show correlation plots for each component of the polarizability tensor of the anthracene crystal, comparing the predicted quantities with the calculated DFPT values over a separate test set of 150 structures (not used for training).
The data shown in Fig. 2 confirms that both methods are very accurate in predicting these tensors, as indeed previous work in the literature has shown [15, 16, 49, 53, 48, 18]. However, we do observe that for the same training dataset size, the MACE- predictions are almost a factor two more accurate than the SA-GPR predictions for diagonal and off-diagonal components of the tensor. One could make the SA-GPR model more accurate by adding more training data, but it is also important to highlight that MACE, as a parametric ML model, does not result in more expensive predictions when increasing the training data set, while SA-GPR, being a non-parametric model, does. This distinction makes MACE the most suitable solution for achieving higher prediction accuracy, especially when applied to larger datasets and system sizes.
In the following, we will use only our two MACE models for the forces and the polarizability predictions for both anthracene and naphthalene.
IV.2 Quality of the Approximations for PO-Raman Signals
Before diving into the analysis of the temperature dependence of the PO-Raman spectra, we analyze the differences between (i) evaluating Eq. 8 with the MACE-MLIP potential for performing the nuclear time evolution combined with the MACE- model for the polarizability evaluations and (ii) evaluating the RGDOS approximation in Eq. 14 with the same MACE-MLIP potential for the nuclear dynamics and fixed Raman tensors computed directly from DFPT. In the following, we name the first procedure “full-ML” and the second “RGDOS-ML”.
In all the calculations performed in this work, the incident and scattered light are polarized in the plane and propagate along the direction in the lab-frame. This means that only the and components of the response are relevant. The PO-Raman spectra we show in the main section are computed in the parallel scattering configuration, as defined in Section II.
The small discrepancies that arise are shown in more detail in Fig. 3c, where the difference between the PO-Raman spectra of anthracene at 100 K obtained with each procedure is shown. Negative values correspond to features that are more intense in full-ML than in RGDOS-ML and vice-versa. In Fig. 3d we report the same difference spectrum, this time for the unpolarized Raman spectrum, as defined in Eq. 2. In general, RGDOS-ML demonstrates remarkable accuracy for PO-Raman analysis in the low-THz region of interest, with discrepancies in the order of of the maximum amplitude. The larger discrepancy is found in the mode of anthracene located below 100 cm-1, where RGDOS-ML overestimates the unpolarized peak amplitude 12%. Consequently, as our focus is limited to first-order Raman – which contains contributions almost solely from the vicinity of -point phonons – we opted to proceed with this framework due to its efficiency and additional advantage of removing uncertainties in the prediction of polarizability tensors.
IV.3 Temperature Dependence of PO-Raman Signals
@@
In Fig. 4 we report the temperature evolution of the parallel configuration PO-Raman spectra of anthracene and naphthalene in the intermolecular motion range of 10 to 170 cm-1, obtained with the RGDOS-ML method. The perpendicular configuration PO-maps can be found in the SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results. (Section S3).Both materials crystallize in the same space group, with two molecules per primitive cell, occupying the same sites in both crystals. Factor group analysis thus leads to the same symmetries for their Raman active modes. Namely, the and modes are respectively the three in-phase and the three out-of-phase librational modes of the two molecular units in the primitive cell, and are Raman active[55, 56]. From the three out-of-phase translational modes, two bear and one bears symmetry, being instead IR active. In our simulations, the crystal orientation is along the crystallographic plane (001), therefore, only the components of the Raman cross-section will be measured.
@@
Focusing on panels a and b of Fig. 4, we can see that for both crystals the peaks corresponding to the modes present a four-fold symmetry in their polarization dependence. On the other hand, modes present a larger variation in their PO patterns. Some modes, like and in anthracene still show a four-fold pattern, but with two maxima of different intensities, while other modes, like in anthracene and in naphthalene show a two-fold oscillation pattern. Also, the mode of naphthalene has a weak PO dependence in the parallel configuration, while the peak has zero intensity in the perpendicular experimental geometry (see Section S3 of SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results.). These observations are intimately connected to the relative magnitudes of the components of the Raman tensors of the respective modes, which will be discussed in detail in Section IV.3.2.

### v2

# Investigating Anharmonicities in Polarization-Orientation Raman Spectra of Acene Crystals with Machine Learning

@@ I Introduction
In this paper, we provide an ab initio quality framework to reproduce, explain and provide quantitative insights into anharmonic effects directly probed by PO-Raman spectroscopy. As a case study, we focus on linear oligoacene molecular crystals, in particular naphthalene and anthracene. The standard approach of combining ab initio molecular dynamics (AIMD) with density functional perturbation theory (DFPT) to obtain Raman spectra guarantees that all orders of anharmonicity of the potential energy surface (PES) are included, but is prohibitively expensive for this task. This is due, for example, to the vast range of experimental conditions one needs to explore and the large cell sizes required to accurately describe the phonon manifold. Furthermore, to capture subtle variations in the angle-resolved Raman intensity, the spectra have to be thoroughly converged, and finite-size artifacts due to periodic boundary conditions need to be mitigated. While ML-based frameworks for computing Raman spectra without explicitly accounting for the crystal orientation have been demonstrated [64, 13, 33, 32], this work instead presents a ML framework to obtain accurate polarization-orientation Raman spectra across multiple temperatures. This extension is non-trivial: PO-Raman intensities provide a significantly more stringent test of ML accuracy than angle-integrated spectra, as subtle errors in polarizability tensor components or normal mode symmetries dramatically alter angular dependence while remaining invisible in unpolarized spectra. We thus combine state-of-the-art equivariant ML models for both interatomic potentials and polarizability tensors with the computationally efficient RGDOS approximation proposed in Refs. [35, 14] to enable quantitative predictions of temperature-dependent anharmonic PO-Raman signals.
In this manuscript, we begin by introducing the theoretical background of PO-Raman scattering, emphasizing the limitations of the harmonic approximation and the need for anharmonic treatments. After benchmarking our ML models and validating the RGDOS approximation against full ML calculations, we simulate the temperature-dependent PO-Raman spectra of anthracene and naphthalene, capturing anharmonic effects. By comparing our results with experimental data, we analyze the polarization dependence of Raman intensities and discuss the subtle deviations from the harmonic predictions. Our findings highlight the interplay between anharmonic lattice dynamics and Raman tensor symmetries, providing insights into the origins of the temperature-dependent behavior of low-frequency Raman modes and suggesting guidelines for the interpretation of the rich amount of information contained in PO-dependent Raman spectra. This work demonstrates the power of combining machine learning with established theoretical frameworks to unravel complex anharmonic phenomena in molecular crystals.
@@ II Theory of Polarization-Orientation Raman Scattering
Single crystal polarization-orientation Raman (also referred to as angle-resolved polarized Raman) experiments are often conducted in a back- (or forward-) scattering geometry, where the incident () and scattered () fields are linearly polarized and propagate through the crystal, which is positioned at a certain orientation [4, 41, 81]. The scattered radiation is then collected with polarization either perpendicular () or parallel () to that of the incident beam. Once the configuration is fixed, the experiment is repeated spanning an angle between and an arbitrary reference, either by rotating the sample or the direction of .
@@
where is the electronic ground state, the summation runs over all the excited electronic states, is the electric dipole operator and denote Cartesian components. With this definition of the polarizability tensor, Eq. 1 becomes the well known Placzek approximation [53, 78] to the full expression of the light scattering cross section derived first by Kramers, Heisenberg and Dirac [46, 26]. Note that, as recently highlighted in Ref. [42], by treating the quantity in Eq. 3 as the polarizability tensor (or dielectric susceptibility) entering Eq. 1, we are already excluding the effect of finite scattering wave-vectors, essentially assuming that only vibrations at the -point of the crystal phonon Brillouin zone are probed in the experiment, i.e. , where the are the propagation directions of incoming and scattered fields. This tends to be a very good assumption for first-order Raman scattering, but breaks down in the context of second-order Raman scattering in polar materials [42].
@@
that will depend trivially on temperature through the population of the initial states. These derivations can also be found in the SM of Ref. [12]. We will often refer to the angular dependence, for a given , as the “PO-pattern” of the corresponding mode in the following.
The result of Eq. 7 is the well known expression for the (harmonic) polarized Raman scattering cross section that is commonly employed to rationalize experimental results in terms of Raman tensors and factor group analysis [24]. As shown above and already clearly stressed in [12], this simple expression is the result of what is often termed a “double harmonic” approximation, as it relies on both harmonic normal modes and the truncation at first order of the expansion in Eq. 4. Clearly, this formula would fall short of accounting for any effect of anharmonic lattice dynamics in the Raman spectrum. Particularly relevant for this work is the fact that in this approximation, the dependence of the (normalized) scattered intensity on the polarization angle is fully determined by the shape, symmetry and relative values of the Raman tensor components of a given mode, and by the geometry of the experiment. Notably, the thermal populations enter only as a multiplicative factor. As the space group of the crystal and the molecular sites do not change with temperature (assuming no phase transitions), the shape and symmetry of are fixed. A common interpretation of this observation is that besides a change in its absolute values, the angular dependence of the intensity calculated with Eq. 7 will be temperature independent in the harmonic picture. As we will see in this paper, though, keeping to Eq. 7 but including temperature-dependent lattice changes can lead to modifications in sign and relative magnitude of certain components of , which in turn lead to pronounced variations in the PO pattern that are not connected to dynamical anharmonic effects.
Beyond this, any further change with temperature of the PO-patterns must then be investigated taking into account the dynamical anharmonicity of molecular vibrations, which presents an added challenge to simulations. In particular, dynamical methods which are not based on perturbation theory for the nuclear degrees of freedom provide an attractive platform to include these effects in their totality. For practical purposes, it is convenient to switch from the energy representation of Eq. 1 to the Heisenberg representation in the time domain, resulting in
@@ III Simulating PO-Raman Signals
Achieving first-principles accuracy in this evaluation relies on being able to obtain calculations of forces and polarizability tensors, through the realization of AIMD trajectories and perturbation theory calculations, resulting in a very high computational cost [36]. Previous studies [63, 62, 60] have used AIMD to successfully compute anharmonic unpolarized or powder Raman spectra of molecular crystals, albeit computing very few trajectories of the order of 10 ps, resulting on a high uncertainty in peak intensities. In the present study, converged peak intensities are necessary in order to capture subtle variations in polarization-orientation patterns at different temperatures.
@@ III.1 Machine Learning Models for Polarizability Tensors
Both methods mentioned above have already been employed in the context of predicting polarizability tensors that were then used to calculate standard vibrational Raman spectra [64, 13, 55, 38, 40, 32]. Here we instead assess them for the more stringent test of predicting the PO-pattern of each Raman peak.
@@ III.2 The RGDOS Approximation to Polarizability Tensors
The RGDOS approximation to the polarizability tensor, as introduced in Ref. [35] and extended for molecular dynamics in Ref. [14], is an efficient approach to reconstruct such tensors of a large supercell as a weighted sum of polarizabilities of the multiple unit cells that constitute it. In Ref. [13], Berger et al. showed the good performance of this approximation in reproducing DFT-quality polarizabilities for large supercells and yielding good Raman spectra, including second-order and resonant contributions. The method relies on the existence of a unique equilibrium reference structure and its corresponding eigenmodes, being therefore unable to describe different phases of a material within a single model, but performing well when no phase transitions occur.
@@
where anharmonic effects are included in the correlation function of the projected supercell displacements. A further possible simplification regards the truncation of Eq. 13 with respect to the number of phonons considered. The evaluation of Eq. 14 then requires fewer Raman tensors and projection coefficients. We demonstrate in the SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results. (Section S2) how retaining only the intermolecular modes in the expansion has no visible effect in the resulting low-frequency spectrum of the polyacene crystals studied here, making this method particularly attractive for the investigation of low THz regions of molecular crystals.
III.3 Proposed Framework
We assessed the quality of the machine-learning interatomic potentials (MLIPs) we trained for the frequencies of the crystal vibrations. A particularly challenging region, where we must guarantee that the errors are acceptable, is the low-THz part of the spectrum, which is determined by weak intermolecular interactions often poorly described in MLIPs that are inherently local. As detailed in Section VI.1, we compared a second-generation Behler-Parrinello high-dimensional neural network potential [10], which we labeled HDNNP-MLIP, and an equivariant graph neural-network MACE [7] interatomic potential, which we labeled MACE-MLIP.
Our benchmarks, detailed in the SM Section S3 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results., show that MACE-MLIP outperforms the HDNNP-MLIP in the prediction of -point harmonic phonon frequencies of the anthracene crystal, resulting in a lower average percentage error of 0.2% against the 1.6% obtained from HDNNP-MLIP. Moreover, MACE-MLIP exhibits a significantly tighter distribution of errors, with a low median and minimal variability, highlighting the consistency and accuracy of this model. In contrast, HDNNP-MLIP shows a broader error distribution, indicating less reliable predictions. When focusing specifically on the low THz region, the error on the harmonic frequency of Raman active modes is almost negligible for the MACE-MLIP model, while the HDNNP-MLIP largely underestimates the harmonic frequency of the three highest frequency librational modes, with errors up to 20%. These conclusions are consistent with the broader benchmark on the accuracy of vibrational properties obtained from MLIPs that we performed in Ref. [33], in which the reasons for the superior performance of MACE-MLIPs for vibrational frequencies of molecular crystals are discussed.
Next, we assessed the quality of our ML models for the prediction of the polarizability tensors. Our benchmarks, detailed in the SM Section S3, show that MACE- predictions are almost a factor two more accurate than the SA-GPR predictions for diagonal and off-diagonal components of the tensor. We conclude that MACE is the most suitable solution for achieving higher prediction accuracy, especially when applied to larger datasets and system sizes.
Lastly, we analyzed the differences between (i) evaluating Eq. 8 with the MACE-MLIP potential for performing the nuclear time evolution combined with the MACE- model for the polarizability evaluations and (ii) evaluating the RGDOS approximation in Eq. 14 with the same MACE-MLIP potential for the nuclear dynamics and fixed Raman tensors computed directly from DFPT. In the following, we name the first procedure “full-ML” and the second “RGDOS-ML”.
In all the calculations performed in this work, the incident and scattered light are polarized in the plane and propagate along the direction in the lab-frame. This means that only the and components of the response are relevant. The PO-Raman spectra we show in the following figures are computed in the parallel scattering configuration, as defined in Section II.
The small discrepancies that arise are shown in more detail in Fig. 1c, where the difference between the PO-Raman spectra of anthracene at 100 K obtained with each procedure is shown. Negative values correspond to features that are more intense in full-ML than in RGDOS-ML and vice-versa. In Fig. 1d we report the same difference spectrum, this time for the unpolarized Raman spectrum, as defined in Eq. 2. In general, RGDOS-ML demonstrates remarkable accuracy for PO-Raman analysis in the low-THz region of interest, with discrepancies in the order of of the maximum amplitude. The larger discrepancy is found in the mode of anthracene located below 100 cm-1, where RGDOS-ML overestimates the unpolarized peak amplitude 12%. Consequently, as our focus is limited to first-order Raman – which contains contributions almost solely from the vicinity of -point phonons – we opted to proceed with the RGDOS-ML framework due to its efficiency and additional advantage of removing uncertainties in the prediction of polarizability tensors.
@@ IV Results and discussion
IV.1 Quasi-Harmonic Approximation to PO-Raman
As a first step towards the analysis of the impact of temperature on the polarization dependence of the Raman intensity, we compute the quasi-harmonic PO-dependence (parallel configuration) of the Raman intensity in anthracene and naphthalene using only the harmonic approximation of Eq. 7. This calculation is purely harmonic and non-dynamical: the Raman tensors are obtained directly through DFPT at the optimized geometries corresponding to different lattice parameters, and no MD simulations are involved. Thermal lattice expansion is taken into account at 100 K and 295 K by using experimentally-determined lattice constants. In this subsection, when we refer to different temperatures, we are simply referring to the corresponding different lattice constants. This approach is commonly referred to as quasi-harmonic approximation.
@@ IV.1 Quasi-Harmonic Approximation to PO-Raman
Both materials studied here crystallize in the same space group, , with two molecules per primitive cell occupying the same sites in both crystals. Factor group analysis thus leads to the same symmetries for their Raman active modes. Namely, the and modes are respectively the three in-phase and the three out-of-phase librational modes of the two molecular units in the primitive cell, and are Raman active [72, 82]. From the three out-of-phase translational modes, two bear and one bears symmetry, being instead IR active. As already mentioned, the crystal orientation in our simulations is along the crystallographic plane (001), meaning only the components of the Raman cross-section will be measured.
Within the quasi-harmonic picture, the polarization pattern, i.e the angular dependence of the Raman intensity for a given mode in Eq. 7, is fundamentally determined by the shape, symmetry, and the relative magnitude and sign of the components of the corresponding . In Section S4 of the SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results. we provide explicit expressions for the general angular dependencies of the Raman intensity patterns generated by and tensors. In the case of naphthalene and anthracene, with the crystals oriented in the crystallographic plane (001), we show how, once the intensity is normalized at each temperature, no variations with lattice expansion can occur in the polarization patterns of modes, both in the parallel and perpendicular configuration, within the harmonic approximation of Eq. 7. The same is true for modes in the perpendicular configuration. In contrast, we show how modes in the parallel configuration are affected by changes in the relative value of their Raman tensor components, which can be induced by thermal lattice expansion.
We report in Fig. 2 the and modes of both crystals as interesting example cases. For completeness, we show the patterns for all modes in parallel and perpendicular configurations in the SM Section S4 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results.. We start by discussing the PO-patterns of the anthracene crystal. All modes show a four-fold oscillating pattern across polarization angles, as exemplified in Fig. 2a, and the modes show more variability (see also Fig. S4 of SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results.). Specifically, and have and tensor components of opposite sign, resulting in a four-fold pattern with two maxima of different intensities, whereas displays a two-fold oscillation due to having these diagonal elements of the same sign. This is an important observation, as despite the fact that a square modulus appears in Eq. 7, the relative sign of the Raman tensor components is qualitatively relevant in the resulting PO-dependence of modes, and can therefore be extracted from a fit of the experimental data. Variations of the PO-patterns at different temperatures are minimal. As a matter of fact, we only observe a change in the relative values of the diagonal components of the mode, resulting in an intensity variation of the relative maximum at 270∘ in Fig. 2a.
@@
In the case of naphthalene, the same analysis holds true for the modes, which show a four-fold oscillating pattern, as expected from the fact that these crystals belong to the same space group. The modes again show either a two-fold or a four-fold pattern with maxima of different intensities, depending on the sign of the diagonal components of their Raman tensors (see Fig. S5 of SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results.). For example, the mode of naphthalene has a weak PO dependence in the parallel configuration, while having zero intensity in the perpendicular one. An inspection of its Raman tensor at 100 K explains this, as the and components are almost equal in magnitude and sign, acting like an identity matrix in Eq. 7.
Quite surprisingly, though, in naphthalene we observe pronounced changes in the overall polarization patterns of modes with thermal lattice expansion. This is particularly evident in the case of mode , where a significant change in the relative values of the diagonal components of the Raman tensor leads to a transition from a four-fold pattern with two maxima of different intensities at 100 K to a two-fold pattern at 295 K (see Fig. 2b). This observation highlights that even within the quasi-harmonic approximation, temperature-induced lattice expansion can lead to substantial modifications in the polarization dependence of certain Raman active modes. In Section S4 of SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results., by comparing the normal modes displacement vectors at the two different temperatures, we show that the and modes of naphthalene present a measurable change in direction with the different lattice parameters, leading to significant variations in their Raman tensor components and consequently in their polarization patterns. This is not the case for anthracene, where the normal modes obtained for the structures at different temperatures overlap almost perfectly with each other.
As a further note, we demonstrate in Section S5 of SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results., that even small inaccuracies in the normal modes can lead to Raman tensors that yield incorrect polarization dependencies. These changes may be overlooked when the variable is integrated out, underlynig again how PO-patterns are a much stricter test for the quality of the theoretical framework, compared to unpolarized Raman spectra.
In summary, once the experimental configuration is fixed (either or ), only the PO-dependence of modes of a certain symmetry can be affected by thermal lattice expansion in the quasi-harmonic approximation. In the case of anthracene and naphthalene, these are the modes in the parallel configuration. The extent to which the PO-pattern of these modes is affected varies significantly between different materials, as exemplified by the comparison between anthracene and naphthalene. While we prove that not all temperature-dependent changes in PO-Raman patterns will stem from dynamical coupling between different vibrational modes, we also provide a clear baseline for the following analysis of temperature-dependent spectra. It is important to note that the changes we capture within the quasi-harmonic approximation in Fig. 2, although sometimes pronounced, are still encoded in the form of second-rank Raman tensors of fixed symmetry. The quasi-harmonic results thus establish a baseline for interpreting temperature effects that arise purely from lattice expansion. Any further changes in the polarization patterns beyond this baseline, as observed in experiment or in the dynamical simulations we present next, must originate from anharmonic dynamical effects involving the vibrational degrees of freedom, where the full, fourth-rank nature of the Raman response function (Eq. 10) is revealed.
IV.2 Temperature Dependence of PO-Raman Signals
We now turn our attention to the investigation of the temperature dependence of the PO-Raman spectra of anthracene and naphthalene crystals. In Fig. 3 we report the temperature evolution of the parallel configuration anharmonic PO-Raman spectra in the intermolecular motion range of 10 to 170 cm-1, obtained with the RGDOS-ML method. The perpendicular configuration PO-spectra can be found in the SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results. (Section S6) for completeness. Computational details on these production simulations are given in Section VI.5.